A drone with a mass of 5,8kg is used to lift a mass of 3,2kg by using an inextensible cable of negligible mass. The tension on a cable is T, and the drone accelerates vertically upwards at 0.9. The lifting force of the drone is 180N.The friction force experienced by the drone and the mass as they move upwards is the same.
1)Calculate the power of the motors of the drone if the drone reaches its maximum constant speed of 19,7
1
Expert's answer
2020-10-07T07:30:35-0400

is the lifting force

is the frictional force

is the mass of the drone

m is the mass of the load

is the acceleration.

Apply Newton's second law and assume that Y-axis looks upward:



"F-mg-Mg-f=(M+m)a,\\
f=F-(m+M)g-(M+m)a=\\
=F-(M+m)(g+a)=\\180-(5.8+3.2)(9.8+0.9)=83.7\text{ N}."